Base of a triangle is 4cm greater than the height. The area is 30cm^2
Find the height and the length

Answer by rajagopalan(174)

Let height=x
then base=x+4
area=1/2*b*h
=(1/2)*(x)*(x+4)
=1/2(x^2+4x)
1/2(x^2+4x)=30
x^2+4x=30*2
x^2+4x-60=0
x^2+10x-6x-60=0
x(x+10)-6(x+10)=0
(x-6)(x+10)=0
giving x=6 or -10
Ignoring negative value
we get x=6
Answer:
Height=6
Base=10
Verification:
Area=1/2*b*h
=1/2*6*10
=3*10
=30 sq units
okay